OpenStudy (anonymous):

Help with my algebra homework. 1.) what is the wuotient in simplified form? state any restrictions on the variable. (x^2-16/x^2+5x+6) / (x^2+5x+4/x^2-2x-8) 2.) simplify the sum (7/a+8) + (7/a^2-64) 3.) simplify the difference (a^2-2a-3/a^2-9a+18) - (a^2-5a-6/a^2+9a+8) 4.) simplify the complex fraction (4/x+3) / ((1/x)+3) 5.) solve the equation. Check the solution (-2/x+4)=(4/x+3) * I have been trying to figure these out all weekend and I can't. Please help.

OpenStudy (campbell_st):

ok... so why you divide by a fraction, you find the reciprocal of the 2nd fraction and multiply so you get \[\frac{x^2 - 16}{x^2 + 5x + 6} \times \frac{x^2 -2x - 8}{x^2 + 5x + 4}\] if you factorise the numerators and denominators the you will see common factors that will cancel \[\frac{(x -4)(x+4)}{(x + 2)(x + 3)} \times \frac{(x -4)(x + 2)}{(x + 1)(x + 4)}\] once you have eliminated the common factors distribute the numerator and denominator for the answer. hope this make sense

OpenStudy (anonymous):

I figured everyone except for one. I cannot get this one (3x-7/x^2) / ((x^2/2)+(2/x)) Can anyone help?

OpenStudy (whpalmer4):

Is that \[\frac{\frac{3x-7}{x^2}}{\frac{x^2}{2}+\frac{2}{x}}\] Yechhhh! Let's simplify the denominator first: \[\frac{x^2}{2} + \frac{2}{x}=\frac{x*x^2}{x*2} + \frac{2*2}{x*2} = \frac{x^3+4}{2x}\] Now remember that dividing fractions can be done by multiplying by the reciprocal giving us \[\frac{3x-7}{x^2}*\frac{2x}{x^3+4} = \frac{2(3x-7)}{x(x^3+4)}\] after we cancel matching terms.

OpenStudy (anonymous):

thanks guys. I sent it in lastnight and got two wrong. Im horrible at algebra

OpenStudy (whpalmer4):

Why don't you show us what you did on the two you got wrong, and we'll tell you where the mistake happened?